OpenStudy (anonymous):

(PLEASE HELP!! It'd be MUCH appreciated!) Graph each equation by finding the vertex and making a x & y table. 1. y = 3/2 |3x-1| 2. y = |1/2x - 4| +4 3. y = |-3x + 2| To find the vertex, you put -b/m for the x coordinate and the y coordinate will be whatever is outside to the right of the absolute value bars. (-b/m, c) I've been trying but for these 3 questions I've been getting weird graphs, they're not making a "v" shape like they're supposed to :(.

OpenStudy (anonymous):

Can you find the vertices?

OpenStudy (anonymous):

Yeah you input 2 numbers that are less than the vertexes that you found and 2 that are greater than it and then you input those numbers into the equation for X. @JakeV8

OpenStudy (anonymous):



OpenStudy (anonymous):

I didn't plot two numbers on either side of the vertex on that one... but you could choose another set. I chose x = 1 first (because it's easy!!) and got y = 3. Notice that it is symmetrical or mirror-imaged, so the point (1,3) "reflects" across the vertex area... it's 2/3 away from (1/3,0), so it reflects on another 2/3 past (1/3,0) to (-1/3,3)

OpenStudy (anonymous):

Another easy one might be x = 0 (These is all for question A, by the way...)

OpenStudy (anonymous):

I can get you the (x,y) tables for each of them how many points do you need? @dellzasaur

OpenStudy (anonymous):

Ok so for A: x = y -5 = 24 -4 = 19.5 -3 = 15 -2 = 10.5 -1 = 6 0 = 1.5 1 = 3 2 = 7.5 3 = 12 4 = 16.5 5 = 21 X-Intercept is 1/3

OpenStudy (anonymous):

@Freyes I need 5 please! :) Thank you!!

OpenStudy (anonymous):

B: x = y -5=10.5 -4=10 -3=9.5 -2=9 -1=8.5 0=8 1=7.5 2=7 3=6.5 4=6 5=5.5 6=5 7=4.5 8=4 9=4.5 10=5

OpenStudy (anonymous):

C: x=y -5=17 -4=14 -3=11 -2=8 -1=5 0=2 1=1 2=4 3=7 4=10 5=13 X-intercept is 2/3

OpenStudy (anonymous):

@Freyes @JakeV8 Omggg thank you guys so much!! I think for one of them a friend said I had to make some new equations? It's either 1 or 2. But thank you both soooo much I really appreciate it!! Life savers!! Haha :)